OpenStudy (anonymous):

What value of x makes the equation true? two-eighths equals twelve over x ninety-six forty-eight eighteen fourteen

OpenStudy (anonymous):

2/8=1/4=0.25

OpenStudy (anonymous):

12/48 = 0.25

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Forty - eight is your answer.

OpenStudy (anonymous):

Yw